### v1

# Colouring ($P_{2}\cup P_{4}$, diamond)-free graphs with $\omega$ colours

Abstract
In this paper, we establish an optimal $\chi$-binding function for $(P_{2}\cup P_{4},\text{ diamond})$-free graphs. We prove that for any graph $G$ in this class, $\chi(G)\leq 4$ when $\omega(G)=2$, $\chi(G)\leq 6$ when $\omega(G)=3$, and $\chi(G)=\omega(G)$ when $\omega(G)\geq 4$, where $\chi(G)$ and $\omega(G)$ denote the chromatic number and clique number of $G$, respectively. This result extends the known chromatic bounds for $(P_{2}\cup P_{3},\text{ diamond})$-free graphs [2, 10, 12] by showing that $(P_{2}\cup P_{4},\text{ diamond})$-free graphs admit the same $\chi$-binding function. It also refines the chromatic bound obtained by Angeliya, Karthick and Huang [1] for $(P_{2}\cup P_{4},\text{ diamond})$-free graphs.
Mathematics Subject Classification: 05C15, 05C17, 05C69, 05C75
Keywords: Graph colouring; $P_{2}\cup P_{4}$-free graphs; Diamond-free graphs.
1 Introduction
All graphs in this paper are finite and simple. We follow [16] for undefined notations and terminology. Let $G$ be a graph with vertex set $V(G)$ and edge set $E(G)$. Two vertices $u,v\in V(G)$ are adjacent if and only if $uv\in E(G)$. We denote by $u\sim v$ if $u,v$ are adjacent and $u\nsim v$ if otherwise. The complement $\overline{G}$ of $G$ is the graph with vertex set $V(G)$ and $uv\in\overline{G}$ if and only if $uv\notin G$.
For a vertex $v\in V(G)$, the neighbourhood of $v$, denoted by $N_{G}(v)$, is the set of all vertices adjacent to $v$. It can be simplified to $N(v)$ when there is no danger of ambiguity. For two disjoint vertex sets $A$ and $B$, we let $[A,B]$ denote the set of edges between $A$ and $B$. We say that $A$ is complete (resp. anticomplete) to $B$ if $|[A,B]|=|A||B|$ (resp. $A,B]=\emptyset$). For a subset $S\subseteq V(G)$, we denote by GSG-S the graph obtained from $G$ by deleting the vertices in $S$ together with their incident edges. We denote the subgraph induced on $X\subseteq V(G)$ by $G[X]$. We will, by a standard abuse of notation, identify a subset $X\subseteq V(G)$ with its induced subgraph $G[X]$ when the context is clear.
A matching in $G$ is a set of edges without common vertices. The vertices incident to the edges of a matching $M$ are saturated by $M$, and a perfect matching is a matching that saturates every vertex of $G$. For bipartite graphs, Hall’s theorem [9] gives a necessary and sufficient condition for the existence of a matching that saturates one part of the bipartition: if $G$ is a bipartite graph with bipartition $(X,Y)$, then $G$ has a matching that saturates $X$ if and only if $|N_{G}(S)|\geq|S|$ for every $S\subseteq X$.
For two graphs $G$ and $H$, we say that $G$ contains $H$ if $H$ is isomorphic to an induced subgraph of $G$. When $G$ does not contain $H$, we say that $G$ is H-free. For a family $\mathcal{H}$ of graphs, we say that G is $\mathcal{H}$-free if $G$ is $H$-free for every $H\in\mathcal{H}$. For two vertex-disjoint graphs $G_{1}$ and $G_{2}$, the union $G_{1}\cup G_{2}$ is the graph with vertex set $V(G_{1}\cup G_{2})=V(G_{1})\cup V(G_{2})$ and edge set $E(G_{1}\cup G_{2})=E(G_{1})\cup E(G_{2})$.
Let $P_{n}$ and $K_{n}$ be a path and a complete graph on $n$ vertices, respectively. The diamond is a graph that consists of a $P_{3}$ with an additional vertex adjacent to all vertices of the $P_{3}$. A hole is an induced cycle of length at least 5, and an antihole is an induced subgraph whose complement graph is a hole. A hole or antihole is odd or even if it is of odd or even length, respectively.
[…]
Figure 1: Illustrations of $P_{2}\cup P_{4}$ and diamond.
We denote by $[k]$ the set $\{1,2,\ldots,k\}$ of the first $k$ positive integers. A $k$-colouring of a graph $G$ is a mapping $c:V\rightarrow[k]$ such that $c(u)\neq c(v)$ whenever $u\sim v$. For a subset $A\subseteq V(G)$, we use $c(A)$ to denote the set of colours assigned to the vertices in $A$ under the colouring $c$. A graph $G$ is $k$-colourable if it admits a $k$-colouring. The chromatic number of $G$, denoted by $\chi(G)$, is the minimum number $k$ for which $G$ is $k$-colourable. We denote by $\omega(G)$ the clique number of $G$ (the cardinality of its largest induced clique), and will simply write $\omega$ when the graph is clear from the context. Clearly, $\chi(G)\geq\omega(G)$. If $\chi(H)=\omega(H)$ for every induced subgraph $H$ of $G$, we say that $G$ is perfect. The famous Strong Perfect Graph Theorem, established by Chudnovsky, Robertson, Seymour and Thomas [6], states that a graph is perfect if and only if it is (odd hole, odd antihole)-free.
The notion of binding functions was introduced by Gyárfás [7] in 1975, extending the concept of perfect graphs. Let $\mathcal{G}$ be a family of graphs. If there exists a real-valued function such that $\chi(G)\leq f(\omega(G))$ holds for every graph $G\in\mathcal{G}$, then we say that $\mathcal{G}$ is $\chi$-bounded, and call $f$ a binding function of $\mathcal{G}$. Gyárfás [7] and Sumner [15] independently presented the following conjecture.
Conjecture $1.$
[7][15] For every forest $T$, the class of $T$-free graphs is $\chi$-bounded.
Gyárfás [8] proved the conjecture for $T=P_{t}$: every $P_{t}$-free graph $G$ satisfied $\chi(G)\leq(t-1)^{\omega(G)-1}$. Note that Gyárfás-Sumner conjecture concerns forests; this naturally leads us to ask what happens when we forbid a disconnected forest, such as $P_{r}\cup P_{s}$.
The problem of colouring $(P_{r}\cup P_{s})$-free graphs (with $r,s\geq 2$) and their subclasses has been studied in various contexts in recent years (see [13] for a comprehensive survey). Currently, the best known binding function for $(P_{2}\cup P_{3})$-free graphs is the same as that for $(P_{2}\cup P_{4})$-free graphs: $f(\omega)=\frac{1}{6}\omega(\omega+1)(\omega+2)$ [2]. Although this function is polynomial, finding a better $\chi$-binding function, preferably linear and optimal, for subclasses of $(P_{2}\cup P_{3})$-free and $(P_{2}\cup P_{4})$-free graphs has received considerable attention.
For example, Char and Karthick [3] proved that $\chi(G)\leq\lceil\frac{5\omega(G)-1}{4}\rceil$ if $G$ is ($P_{2}\cup P_{3}$, gem)-free with $\omega(G)\geq 4$. Li, Li and Wu [11] showed that if a graph $G$ is ($P_{2}\cup P_{3}$, house)-free with $\omega(G)\geq 2$, then it satisfies the optimal bound $\chi(G)\leq 2\omega(G)$. Note that the class of ($P_{r}\cup P_{s}$, diamond)-free graphs is a subclass of ($P_{r+s+1}$, diamond)-free graphs. For the latter, Schiermeyer and Randerath [13] proved the following general result:
Lemma 1.2.
[13] Let $G$ be a connected ($P_{t}$, diamond)-free graph for $t\geq 4$. Then $\chi(G)\leq(t-2)(\omega(G)-1)$.
By Lemma 1.2, every ($P_{r}\cup P_{s}$, diamond)-free graph satisfies $\chi(G)\leq(r+s-1)(\omega(G)-1)$. This leaves room for refinement. For instance, Bharathi and Choudum [2] established that for the class of ($P_{2}\cup P_{3}$, diamond)-free graphs, $\chi(G)\leq 4$ when $\omega(G)=2$, a bound which is tight, and that the graph is perfect for $\omega(G)\geq 5$. For the same class, Karthick and Mishra [10] showed that $\chi(G)\leq 6$ when $\omega(G)=3$, which is also optimal. Furthermore, Prashant, Francis and Raj [12] proved that $\chi(G)=4$ for $\omega(G)=4$.
As a superclass of ($P_{2}\cup P_{3}$, diamond)-free graphs, the class of ($P_{2}\cup P_{4}$, diamond)-free graphs was studied by Chen and Zhang [4]. They established the following chromatic bounds: $\chi(G)\leq 4$ for $\omega(G)=2$, $\chi(G)\leq 7$ for $\omega(G)=3$, $\chi(G)\leq 9$ for $\omega(G)=4$, and $\chi(G)\leq 2\omega(G)+1$ for $\omega(G)\geq 5$. In [4], they also derived binding functions for ($P_{2}\cup P_{4}$, gem)-free and ($P_{2}\cup P_{4}$, butterfly)-free graphs, and proved that ($P_{2}\cup P_{4}$, diamond, $C_{5}$)-free graphs are perfect when $\omega(G)\geq 5$. More recently, Angeliya, Karthick and Huang [1] made significant progress by demonstrating that for ($P_{2}\cup P_{4}$, diamond)-free graphs with $\omega(G)\geq 3$, $\chi(G)\leq\max\{6,\omega(G)\}$. Their work also settled the case for $\omega(G)=4$ by proving $\chi(G)=4$. However, the bound was not tight for $\omega(G)=5$, leaving room for further refinement. Moreover, they improved on a result from [4] by showing that if $G$ is ($P_{2}\cup P_{4}$, diamond, $C_{5}$)-free with $\omega(G)\geq 4$, then $G$ is perfect. Subsequently, Chen, Wu and Zhang [5] showed that ($P_{2}\cup P_{4}$, bull, diamond)-free graphs satisfy $\chi(G)\leq\max\{4,\omega(G)\}$. In the same work, they also established a decomposition theorem for $(P_{2}\cup P_{4})$-free graphs with $\omega(G)\leq 2$, showing that such graphs have $\chi(G)\leq 4$ and that those attaining $\chi(G)=4$ contain the Mycielski–Grötzsch graph.
1.1 Our Contributions
This paper establishes a complete characterisation of the $\chi$-binding function for the class of ($P_{2}\cup P_{4}$, diamond)-free graphs. Our result brings two improvements. First, it shows that $(P_{2}\cup P_{4},\text{ diamond})$-free graphs admit the same $\chi$-binding function as $(P_{2}\cup P_{3},\text{ diamond})$-free graphs [2, 10, 12]. Second, it refines the chromatic bound for $(P_{2}\cup P_{4},\text{ diamond})$-free graphs that was derived by Angeliya, Karthick and Huang [1].
Through a different structural decomposition approach, we obtain the optimal $\chi$-binding function across all values of $\omega(G)$, thereby providing a complete solution to this colouring problem.
Theorem 1.3.
Let $G$ be a ($P_{2}\cup P_{4}$, diamond)-free graph. Then
$\chi(G)\begin{cases}\leq 4,&\text{if }\omega(G)=2,\\ \leq 6,&\text{if }\omega(G)=3,\\ =\omega(G),&\text{if }\omega(G)\geq 4.\end{cases}$
Furthermore, each of these bounds is tight.
It is important to note that, however, a ($P_{2}\cup P_{4}$, diamond)-free graph with clique number at least four is not necessarily perfect. To see this, consider the graph $H_{n}$ ($n\geq 4$) formed by identifying an edge of a $K_{n}$ with an edge of a $C_{5}$ (Figure 2 illustrates this construction for $n=5$). One can verify that $H_{n}$ is ($P_{2}\cup P_{4}$, diamond)-free. Clearly, $\chi(H_{n})=\omega(H_{n})=n$. By the Strong Perfect Graph Theorem [6], $H_{n}$ is imperfect since it contains an induced $C_{5}$.
Figure 2: An imperfect ($P_{2}\cup P_{4}$, diamond)-free graph $G$ with $\chi(G)=\omega(G)=5$.
The remainder of this paper is organised as follows. Section 2 presents fundamental structural properties of ($P_{2}\cup P_{4}$, diamond)-free graphs. This structural analysis leads to two primary cases, which are treated in detail in Sections 3 and 4, respectively. By synthesising these results we prove our main theorem in Section 5.
We conclude this section with the following lemma, which will be used several times in the sequel.
Lemma 1.4.
Let $G$ be a graph containing two disjoint cliques $X$ and $Y$ such that $\min\{|X|,|Y|\}\geq 2$ and $\max\{|X|,|Y|\}\geq 3$. If $[X,Y]$ induces a nonempty matching, then for any $x\in X$ and $y\in Y$, $G[X\cup Y]$ contains a $P_{4}$ whose vertex set includes both $x$ and $y$.
Proof.
By symmetry, assume that $|X|\geq 3$. Let $x\in X$ and $y\in Y$. We show that $G[X\cup Y]$ contains a $P_{4}$ whose vertex set contains both $x$ and $y$.
First, consider the case when $x$ is anticomplete to $Y$. If $y$ is also anticomplete to $X$, then there exist vertices $x^{\prime}\in X$ and $y^{\prime}\in Y$ such that $x^{\prime}\sim y^{\prime}$. Hence, $\{x,x^{\prime},y^{\prime},y\}$ induces a $P_{4}$. If $y$ is not anticomplete to $X$, then there exists $x^{\prime}\in X$ such that $y\sim x^{\prime}$. In this case, $\{x,x^{\prime},y,y^{\prime\prime}\}$ induces a $P_{4}$, where $y^{\prime\prime}$ is any vertex in $Y\setminus\{y\}$.
Next, consider the case when $x$ is not anticomplete to $Y$. If $x\sim y$, then there exist vertices $x^{\prime}\in X$ and $y^{\prime}\in Y$ such that $x^{\prime}\nsim y^{\prime}$, since $|X|\geq 3$. Hence, $\{x^{\prime},x,y,y^{\prime}\}$ induces a $P_{4}$. If $x\nsim y$, then there exists $y^{\prime}\in Y$ such that $x\sim y^{\prime}$. Since $|X|\geq 3$, we can choose $x^{\prime}\in X$ such that $y\nsim x^{\prime}$. Thus, $\{x^{\prime},x,y^{\prime},y\}$ induces a $P_{4}$. ∎
2 The Structure of ($P_{2}\cup P_{4}$, diamond)-Free Graphs
Let $G$ be a $(P_{2}\cup P_{4},\text{ diamond})$-free graph with $\omega(G)\geq 2$. Let $A=\{a_{1},a_{2},\ldots,a_{\omega}\}$ be a maximum clique in $G$, and let H=GAH=G-A, which we assume is nonempty. For every $v\in V(H)$ we have $|N_{A}(v)|\leq 1$: if $\omega\geq 3$ this follows from the fact that $G$ is diamond-free, and the case $\omega=2$ is trivial. We now partition $V(H)$ into the following subsets, where $i\in[\omega]$:
$C_{i}$ $=\{v\in V(H)\mid N_{A}(v)=\{a_{i}\}\},$
$C_{0}$ $=\{v\in V(H)\mid N_{A}(v)=\emptyset\}.$
Lemma 2.1.
For all $1\leq i\neq j\leq\omega$, every subgraph induced by a subset of $V(H)\setminus(C_{i}\cup C_{j})$ is perfect.
Proof.
This follows directly from the facts that $G$ is $(P_{2}\cup P_{4})$-free and that every $P_{4}$-free graph is perfect [14]. ∎
For the case $\omega=2$, the following lemma shows that $\chi(G)\leq 4$.
Lemma 2.2.
If $G$ is a ($P_{2}\cup P_{4}$, diamond)-free graph with $\omega=2$, then $\chi(G)\leq 4$.
Proof.
Since $\omega=2$, both $C_{1}$ and $C_{2}$ are stable. By Lemma 2.1, $G[C_{0}]$ is perfect, and hence it can be coloured with at most two colours. Now we may colour $a_{1}$ and $a_{2}$ with $1$ and $2$, respectively, and colour $C_{1}$ and $C_{2}$ with $2$ and $1$, respectively; finally colour $C_{0}$ with at most two colours from $\{3,4\}$. ∎
By Lemma 2.2, we may assume that $\omega(G)\geq 3$ throughout this paper.
Let $B=\{b_{1},b_{2},\ldots,b_{k}\}$ be a maximum clique in $H$, where $1\leq\omega(B)=k\leq\omega$. For $\omega\geq 4$, the next lemma provides an upper bound on the chromatic number of $G$ in terms of $\omega$ and $k$.
Lemma 2.3.
If $G$ is a ($P_{2}\cup P_{4}$, diamond)-free graph with $\omega\geq 4$, then $\chi(G)\leq\max\{2k,\omega\}$.
Proof.
Let
$X=C_{0}\cup C_{1}\cup\cdots\cup C_{\lfloor\frac{\omega}{2}\rfloor}\quad\text{and}\quad Y=C_{\lfloor\frac{\omega}{2}\rfloor+1}\cup\cdots\cup C_{\omega}.$
By Lemma 2.1, both $G[X]$ and $G[Y]$ are perfect, and hence each of $X$ and $Y$ can be coloured with at most $k$ colours. Moreover, $V(G)=A\cup X\cup Y$.
Now we may colour $a_{i}$ with $i$, and colour $X$ with colours from $\{\lfloor\frac{\omega}{2}\rfloor+1,\lfloor\frac{\omega}{2}\rfloor+2,\ldots,\lfloor\frac{\omega}{2}\rfloor+k\}$. If $k\leq\lfloor\frac{\omega}{2}\rfloor$, then we colour $Y$ with colours from $\{1,2,\ldots,\lfloor\frac{\omega}{2}\rfloor\}$. If $k>\lfloor\frac{\omega}{2}\rfloor$, then we colour $Y$ with colours from $\{1,2,\ldots,\lfloor\frac{\omega}{2}\rfloor,\lfloor\frac{\omega}{2}\rfloor+k+1,\ldots,2k\}$. In both cases we obtain $\chi(G)\leq\max\{2k,\omega\}$. ∎
For $\omega=3$ and $k\leq 2$, next lemma shows that $\chi(G)\leq 6$.
Lemma 2.4.
If $G$ is a ($P_{2}\cup P_{4}$, diamond)-free graph with $\omega=3$ and $k\leq 2$, then $\chi(G)\leq 6$.
Proof.
By Lemma 2.1, each of the sets $C_{0}\cup C_{1}$, $C_{2}$, and $C_{3}$ induces a perfect graph. Since $k\leq 2$, each of them can be coloured with at most two colours. Hence, we may colour $a_{i}$ with colour $i$ for $i\in[3]$, colour $C_{0}\cup C_{1}$ with colours from $\{2,3\}$, colour $C_{2}$ with colours from $\{1,4\}$, and colour $C_{3}$ with colours from $\{5,6\}$. ∎
By Lemmas 2.3 and 2.4, if $\omega=3$ and $k\leq 2$, then $\chi(G)\leq 6$; whereas if $\omega\geq 4$ and $k\leq 2$, then $\chi(G)=\omega$. Hence, for the remainder of this paper we assume that $k\geq 3$.
Since $H$ is diamond-free and $B$ is a maximum clique of $H$, and $k\geq 3$, for every $v\in V(H)\setminus B$ we have $|N_{B}(v)|\leq 1$. We partition $V(H)\setminus B$ into the following subsets, where $i\in[\omega]$ and $j\in[k]$:
$R_{j}$ $=\{v\in V(H)\setminus B\mid N_{A}(v)=\emptyset,N_{B}(v)=\{b_{j}\}\},$
$S_{i}^{j}$ $=\{v\in V(H)\setminus B\mid N_{A}(v)=\{a_{i}\},N_{B}(v)=\{b_{j}\}\},$
$T_{i}$ $=\{v\in V(H)\setminus B\mid N_{B}(v)=\emptyset,N_{A}(v)=\{a_{i}\}\},$
$Z$ $=\{v\in V(H)\setminus B\mid N_{A}(v)=\emptyset,N_{B}(v)=\emptyset\}.$
Since $G$ is diamond-free, each $S_{i}^{j}$ is stable. Note that $\omega(R_{j}),\omega(T_{i})\leq\omega-1$ and $\omega(Z)\leq k$. Furthermore, by Lemma 2.1, each of the sets $R_{j}$, $T_{i}$, and $Z$ induces a perfect graph. We then define
$R=\bigcup_{j=1}^{k}R_{j},\quad S=\bigcup_{j=1}^{k}\bigcup_{i=1}^{\omega}S_{i}^{j},\quad T=\bigcup_{i=1}^{\omega}T_{i}.$
Note that
$V(G)=A\cup B\cup R\cup S\cup T\cup Z.$
If we define
$C_{j}^{\prime}$ $=\{v\in V(H)\setminus B\mid N_{B}(v)=\{b_{j}\}\},$
$C_{0}^{\prime}$ $=\{v\in V(H)\setminus B\mid N_{B}(v)=\emptyset\},$
then Table 1 illustrates the relations among the sets in the partition introduced earlier.
$\cap$ $C_{0}^{\prime}$ $C_{1}^{\prime}$ $C_{2}^{\prime}$ $C_{3}^{\prime}$ $\cdots$ $C_{k}^{\prime}$
$C_{0}$ $Z$ $R_{1}$ $R_{2}$ $R_{3}$ $\cdots$ $R_{k}$
$C_{1}$ $T_{1}$ $S_{1}^{1}$ $S_{1}^{2}$ $S_{1}^{3}$ $\cdots$ $S_{1}^{k}$
$C_{2}$ $T_{2}$ $S_{2}^{1}$ $\ddots$ $\vdots$
$C_{3}$ $T_{3}$ $S_{3}^{1}$ $\ddots$ $\vdots$
$\vdots$ $\vdots$ $\vdots$ $\ddots$ $\vdots$
$C_{\omega}$ $T_{\omega}$ $S_{\omega}^{1}$ $\cdots$ $\cdots$ $\cdots$ $S_{\omega}^{k}$
Table 1: Partitions of $V(H)\setminus B$
With this partition, we now conclude Section 2 by presenting the following two claims.
Claim 2.1.
If there exists a vertex $v\in A$ such that $|N_{B}(v)|\geq 2$, then $\chi(G)=\omega$.
Proof.
Without loss of generality, let $a_{1}\in A$ be a vertex such that $|N_{B}(a_{1})|\geq 2$. Since $G$ is diamond-free, $a_{1}$ is complete to $B$; that is, $B\subseteq C_{1}$. Thus, $\{a_{2},a_{3}\ldots,a_{\omega}\}$ is anticomplete to $\{b_{1},b_{2},\ldots,b_{k}\}$. It follows that $k\leq\omega-1$. Otherwise, $\{a_{1}\}\cup B$ would be a clique on $\omega+1$ vertices, a contradiction. Recall that we have already assumed $k\geq 3$ in the previous discussion; hence, $\omega\geq 4$. We consider the following partition of $G$:
$V(G)=A\cup B\cup R\cup S\cup(T\setminus T_{1})\cup(Z\cup T_{1}).$
Next, we list several properties of the sets $R$, $S$, $(T\setminus T_{1})$, and $(Z\cup T_{1})$.
(1) $S_{1}^{j}=\emptyset$ for all $j\in[k]$.
This follows directly from the fact that $G$ is diamond-free.
(2) At most one set among $S_{i}^{1},S_{i}^{2},\ldots,S_{i}^{k}$ is nonempty, and at most one set among $S_{2}^{j},S_{3}^{j},\ldots,S_{\omega}^{j}$ is nonempty, for all $i\in[\omega]\setminus\{1\}$ and $j\in[k]$.
The two cases admit similar proofs; we therefore prove only the first, that is, at most one set among $S_{i}^{1},S_{i}^{2},\ldots,S_{i}^{k}$ is nonempty. To the contrary, suppose without loss of generality that there exist $x\in S_{2}^{1}$ and $y\in S_{2}^{2}$. If $x\sim y$, then $\{a_{3},a_{4},b_{3},b_{1},x,y\}$ induces a $P_{2}\cup P_{4}$. If $x\nsim y$, then $\{a_{3},a_{4},x,b_{1},b_{2},y\}$ induces a $P_{2}\cup P_{4}$. Both cases lead to a contradiction.
(3) $Z\cup T_{1}$ is anticomplete to $R\cup S\cup(T\setminus T_{1})$.
Suppose, to the contrary, that there exist vertices $x\in Z\cup T_{1}$ and $y\in R\cup S\cup(T\setminus T_{1})$ such that $x\sim y$. If $y\in R\cup S$, then, by symmetry, we may assume that $y\sim b_{1}$. Since $\omega\geq 4$, there exist $a_{i},a_{j}\in A\setminus\{a_{1}\}$ that are anticomplete to $x$ and $y$. Thus, $\{a_{i},a_{j},x,y,b_{1},b_{2}\}$ induces a $P_{2}\cup P_{4}$. If $y\in T\setminus T_{1}$, then, by symmetry, we may assume that $y\sim a_{2}$. Since $k\geq 3$, there exist $b_{p},b_{q}\in B$ that are anticomplete to $x$ and $y$. Hence, $\{b_{p},b_{q},x,y,a_{2},a_{3}\}$ induces a $P_{2}\cup P_{4}$. In both cases, we obtain a contradiction.
(4) $R$ is anticomplete to $S\cup(T\setminus T_{1})\cup(Z\cup T_{1})$.
By (3) and symmetry, it suffices to show that $R_{1}$ is anticomplete to $S\cup(T\setminus T_{1})$. Suppose to the contrary that there exist vertices $x\in R_{1}$ and $y\in S\cup(T\setminus T_{1})$ such that $x\sim y$. For the case $y\in S$, if $y\nsim b_{1}$, then since $k\geq 3$, there exists $b_{t}\in B\setminus\{b_{1}\}$ nonadjacent to $y$; and since $\omega\geq 4$, there exist $a_{i},a_{j}\in A\setminus\{a_{1}\}$ anticomplete to $x$ and $y$. Thus, $\{a_{i},a_{j},b_{t},b_{1},x,y\}$ induces a $P_{2}\cup P_{4}$. If $y\sim b_{1}$ and $y\sim a_{p}$, then $b_{2}$ and $b_{3}$ are anticomplete to $x$ and $y$. Hence, $\{b_{2},b_{3},x,y,a_{p},a_{q}\}$ induces a $P_{2}\cup P_{4}$, where $a_{q}\in A\setminus\{a_{1},a_{p}\}$. For the case $y\in T\setminus T_{1}$, since $\omega\geq 4$, there exist $a_{i},a_{j}\in A\setminus\{a_{1}\}$ that are anticomplete to $x$ and $y$. Thus, $\{a_{i},a_{j},b_{2},b_{1},x,y\}$ induces a $P_{2}\cup P_{4}$. All cases lead to a contradiction.
(5) $R_{i}$ is anticomplete to $R_{j}$ for distinct $i,j\in[k]$.
For the sake of contradiction, suppose by symmetry that there exist $x\in R_{1}$ and $y\in R_{2}$ such that $x\sim y$. Note that $\{a_{2},a_{3},b_{3},b_{1},x,y\}$ induces a $P_{2}\cup P_{4}$, a contradiction.
(6) $T\setminus T_{1}$ is anticomplete to $R\cup S\cup(Z\cup T_{1})$
(7) $T_{i}$ is anticomplete to $T_{j}$ for distinct $i,j\in[\omega]\setminus\{1\}$.
The proofs of (6) and (7) are analogous to those of (4) and (5), and hence are omitted.
(8) $S$ is anticomplete to $R\cup(T\setminus T_{1})\cup(Z\cup T_{1})$.
This follows from (3), (4), and (6).
By (1) and (2) and by symmetry, we may let $S=\bigcup_{i=1}^{k}S_{i+1}^{i}$. By Lemma 2.1, each of the sets $R_{j}$, $T_{i}$, and $Z\cup T_{1}$ induces a perfect graph. Recall that $k\leq\omega-1$. Hence, each of these sets can be coloured with at most $\omega-1$ colours. Finally, we colour $G$ with $\omega$ colours as follows.
  • Colour $a_{i}$ with colour $i$ for $i\in[\omega]$.
  • Colour $b_{i}$ with colour $i+1$ for $i\in[k]$.
  • By (8), colour $S_{i+1}^{i}$ with colour $i$ for $i\in[k]$ (recall that $S_{i+1}^{i}$ is stable).
  • By (3), colour $Z\cup T_{1}$ with at most $\omega-1$ colours from $\omega]\setminus\{1\}$.
  • By (4) and (5), colour $R_{i}$ with at most $\omega-1$ colours from $\omega]\setminus\{i+1\}$ for $i\in[k]$.
  • By (6) and (7), colour $T_{i}$ with at most $\omega-1$ colours from $\omega]\setminus\{i\}$ for $i\in[\omega]\setminus\{1\}$.
This completes the proof of Claim 2.1. ∎
By Claim 2.1, we may assume that for every vertex $v\in A$, $|N_{B}(v)|\leq 1$. Thus,
$\{a_{1},a_{2},\ldots,a_{\omega}\},\{b_{1},b_{2},\ldots,b_{k}\}]$ is a matching in $G$. ($\ast$)
Claim 2.2.
If $H$ contains a $K_{k}$ whose vertex set is anticomplete to $A$, then $\chi(G)=\omega$.
Proof.
Let $B$ be the $K_{k}$ that is anticomplete to $A$. We first list some properties of $Z$, $R$, $S$, and $T$. The proofs of these properties are analogous to those given in the proof of Claim 2.1, and hence are omitted.
(1) At most one set among $S_{i}^{1},S_{i}^{2},\ldots,S_{i}^{k}$ is nonempty, and at most one set among $S_{1}^{j},S_{2}^{j},\ldots,S_{\omega}^{j}$ is nonempty, for all $i\in[\omega]$ and $j\in[k]$.
(2) $Z$ is anticomplete to $R\cup S\cup T$.
(3) $R$ is anticomplete to $S\cup T\cup Z$.
(4) $R_{i}$ is anticomplete to $R_{j}$ for distinct $i,j\in[k]$.
(5) $T$ is anticomplete to $R\cup S\cup Z$
(6) $T_{i}$ is anticomplete to $T_{j}$ for distinct $i,j\in[\omega]$.
(7) $S$ is anticomplete to $R\cup T\cup Z$.
By (1) and by symmetry, we may let $S=\bigcup_{i=1}^{k}S_{i}^{i}$. By Lemma 2.1, each of the sets $R_{j}$, $T_{i}$, and $Z$ induces a perfect graph. Since $\omega(R_{j}),\omega(T_{i})\leq\omega-1$ and $\omega(Z)\leq k\leq\omega$, we may colour them with at most $\omega-1$, $\omega-1$, and $\omega$ colours, respectively. We now present an $\omega$-colouring of $G$ as follows.
  • Colour $a_{i}$ with colour $i$ for $i\in[\omega]$.
  • Colour $b_{i}$ with colour $i$ for $i\in[k]$.
  • By (7), if $k=1$, then colour $S_{1}^{1}$ with colour $2$. If $k\geq 2$, then colour $S_{i}^{i}$ with colour $i+1$ for $i\in[k-1]$, and colour $S_{k}^{k}$ with colour 1.
  • By (2), colour $Z$ with at most $\omega$ colours from $\omega]$.
  • By (3) and (4), colour $R_{i}$ with at most $\omega-1$ colours from $\omega]\setminus\{i\}$ for $i\in[k]$.
  • By (5) and (6), colour $T_{i}$ with at most $\omega-1$ colours from $\omega]\setminus\{i\}$ for $i\in[\omega]$.
This completes the proof of Claim 2.2. ∎
By Claim 2.2, we may assume that
$H$ contains no $K_{k}$ whose vertex set is anticomplete to $A$. ($\ast\ast$)
3 The Case $k=3$
We continue with the notation and definitions introduced earlier. By Lemma 2.3, if $\omega\geq 6$ and $k=3$, then $\chi(G)=\omega$. Hence, throughout this section we may assume that $\omega\in\{3,4,5\}$. The aim of this section is to prove the following three lemmas.
Lemma 3.1.
If $G$ is a $(P_{2}\cup P_{4},\text{ diamond})$-free graph with $\omega=3$ and $k=3$, then $\chi(G)\leq 6$.
Lemma 3.2.
If $G$ is a $(P_{2}\cup P_{4},\text{ diamond})$-free graph with $\omega=4$ and $k=3$, then $\chi(G)=4$.
Lemma 3.3.
If $G$ is a $(P_{2}\cup P_{4},\text{ diamond})$-free graph with $\omega=5$ and $k=3$, then $\chi(G)=5$.
Proof of Lemma 3.1.
We first claim that if there exists $C_{i}$ among $C_{1},C_{2},C_{3}$ such that $\omega(C_{0}\cup C_{i})\leq 2$, then $\chi(G)\leq 6$. Without loss of generality, assume that $\omega(C_{0}\cup C_{1})\leq 2$. Note that $\omega(C_{2}),\omega(C_{3})\leq 2$, and by Lemma 2.1, each of $C_{0}\cup C_{1}$, $C_{2}$, and $C_{3}$ induces a perfect graph. Thus, we may use at most two colours to colour each of $C_{0}\cup C_{1}$, $C_{2}$, and $C_{3}$. Then we colour $a_{1},a_{2},a_{3}$ with colours 1, 2, and 3, respectively; colour $C_{0}\cup C_{1}$ with colours from $\{2,3\}$, colour $C_{2}$ with colours from $\{1,4\}$, and colour $C_{3}$ with colours from $\{5,6\}$.
Hence, we may assume that for each $C_{i}$ among $C_{1},C_{2},$ and $C_{3}$, the subgraph $G[C_{0}\cup C_{i}]$ contains a $K_{3}$. Since $k=3$, we may regard $B$ as such a $K_{3}$. Then, by ($\ast$2), any $K_{3}$ in $G[C_{0}\cup C_{i}]$ has at most one vertex in $C_{i}$. Moreover, by ($\ast\ast$2), $G[C_{0}]$ is $K_{3}$-free. Therefore, every $K_{3}$ contained in $G[C_{0}\cup C_{i}]$ consists of exactly one vertex in $C_{i}$ and two vertices in $C_{0}$.
Let $X\subseteq C_{0}\cup C_{1}$ and $Y\subseteq C_{0}\cup C_{2}$ be two vertex sets such that both $X$ and $Y$ induce a $K_{3}$. Let $X=\{x_{0},x_{0}^{\prime},x_{1}\}$ and $Y=\{y_{0},y_{0}^{\prime},y_{2}\}$, where $\{x_{0},x_{0}^{\prime},y_{0},y_{0}^{\prime}\}\subseteq C_{0}$, $x_{1}\in C_{1}$, and $y_{2}\in C_{2}$. Since $k=3$ and $G$ is diamond-free, $|X\cap Y|\leq 1$.
If $|X\cap Y|=1$, then we may assume that $x_{0}=y_{0}$ (see Figure 3(a)). Since $H$ is both $K_{4}$-free and diamond-free, $\{x_{0}^{\prime},x_{1}\}$ is anticomplete to $\{y_{0}^{\prime},y_{2}\}$. Hence, $\{x_{0}^{\prime},x_{1},y_{0}^{\prime},y_{2},a_{2},a_{3}\}$ induces a $P_{2}\cup P_{4}$, a contradiction.
If $|X\cap Y|=0$ (see Figure 3(b)), then since $k=3$ and $G$ is diamond-free, $\{x_{0},x_{0}^{\prime}\},Y]$ is a matching in $G$. If $\{x_{0},x_{0}^{\prime}\}$ is anticomplete to $Y$, then $\{x_{0},x_{0}^{\prime},y_{0},y_{2},a_{2},a_{3}\}$ induces a $P_{2}\cup P_{4}$, a contradiction. Hence, the matching $\{x_{0},x_{0}^{\prime}\},Y]$ is nonempty. By Lemma 1.4, $G[\{x_{0},x_{0}^{\prime}\}\cup Y]$ contains a $P_{4}$; the vertices of this $P_{4}$, together with $\{a_{1},a_{3}\}$, induce a $P_{2}\cup P_{4}$, a contradiction.
Thus, at least one of $G[C_{0}\cup C_{1}]$ and $G[C_{0}\cup C_{2}]$ is $K_{3}$-free, and hence $\chi(G)\leq 6$. ∎
$x_{0}^{\prime}$$x_{0}$$y_{0}^{\prime}$$x_{1}$$y_{2}$$a_{1}$$a_{2}$$a_{3}$
(a) The case $|X\cap Y|=1$.
$x_{0}^{\prime}$$x_{0}$$y_{0}$$y_{0}^{\prime}$$x_{1}$$y_{2}$$a_{1}$$a_{2}$$a_{3}$
(b) The case $|X\cap Y|=0$.
Figure 3: Two cases in the proof of Lemma 3.1.
Proof of Lemma 3.2.
We first claim that if there exist $C_{i},C_{j}$ among $C_{1}$, $C_{2}$, $C_{3}$, and $C_{4}$ such that $\omega(C_{0}\cup C_{i}\cup C_{j})\leq 2$, then $\chi(G)=4$. Without loss of generality, assume that $\omega(C_{0}\cup C_{1}\cup C_{2})\leq 2$. Since $k=3$, if $G[C_{3}\cup C_{4}]$ contains a $K_{3}$, then we may take this $K_{3}$ as $B$, which contradicts ($\ast$2). Hence, $G[C_{3}\cup C_{4}]$ is $K_{3}$-free, and thus $\omega(C_{3}\cup C_{4})\leq 2$. We may then colour $a_{i}$ with colour $i$ for $i\in[4]$, colour $C_{0}\cup C_{1}\cup C_{2}$ with colours from $\{3,4\}$, and colour $C_{3}\cup C_{4}$ with colours from $\{1,2\}$.
Thus, we may assume that for any $C_{i},C_{j}$ among $C_{1}$, $C_{2}$, $C_{3}$, and $C_{4}$, the subgraph $G[C_{0}\cup C_{i}\cup C_{j}]$ contains a $K_{3}$. Since $k=3$, we may regard $B$ as such a $K_{3}$. By ($\ast$2), any $K_{3}$ in $G[C_{0}\cup C_{i}\cup C_{j}]$ has at most one vertex in $C_{i}$ and at most one vertex in $C_{j}$. Moreover, by ($\ast\ast$2), the subgraph $G[C_{0}]$ is $K_{3}$-free. Therefore, every $K_{3}$ contained in $G[C_{0}\cup C_{i}\cup C_{j}]$ contains at least one vertex in $C_{0}$ and at most two vertices in $C_{0}$. Let $X\subseteq C_{0}\cup C_{1}\cup C_{2}$ be a vertex set such that $G[X]$ is a $K_{3}$.
If $|X\cap C_{0}|=2$, then by symmetry we may assume that $|X\cap C_{1}|=1$ (the case $|X\cap C_{2}|=1$ is analogous). Let $X=\{x_{0},x_{0}^{\prime},x_{1}\}$, where $x_{0},x_{0}^{\prime}\in C_{0}$ and $x_{1}\in C_{1}$. Now consider a vertex set $Y\subseteq C_{0}\cup C_{3}\cup C_{4}$ that induces a $K_{3}$. By symmetry, we may assume that $|Y\cap C_{4}|=1$ (the case $|Y\cap C_{3}|=1$ is analogous). Let $Y=\{y_{0},y_{4},y\}$, where $y_{0}\in C_{0}$, $y_{4}\in C_{4}$, and $y\in C_{0}$ or $C_{3}$. Since $k=3$ and $G$ is diamond-free, we have $|X\cap Y|\leq 1$.
For the case $|X\cap Y|=1$, we may assume that $x_{0}=y_{0}$ (see Figure 4(a)). Since $H$ is $K_{4}$-free and diamond-free, $\{x_{0}^{\prime},x_{1}\}$ is anticomplete to $\{y,y_{4}\}$. Thus, $\{x_{0}^{\prime},x_{1},y,y_{4},a_{4},a_{2}\}$ induces a $P_{2}\cup P_{4}$, a contradiction. For the case $|X\cap Y|=0$ (see Figure 4(b)), since $k=3$ and $G$ is diamond-free, $\{x_{0},x_{0}^{\prime}\},Y]$ is a matching in $G$. If $\{x_{0},x_{0}^{\prime}\}$ is anticomplete to $Y$, then $\{x_{0},x_{0}^{\prime},y_{0},y_{4},a_{4},a_{2}\}$ induces a $P_{2}\cup P_{4}$, a contradiction. Hence, $\{x_{0},x_{0}^{\prime}\},Y]$ is a nonempty matching in $G$. By Lemma 1.4, $G[\{x_{0},x_{0}^{\prime}\}\cup Y]$ contains a $P_{4}$. The vertices of this $P_{4}$, together with $\{a_{1},a_{2}\}$, induce a $P_{2}\cup P_{4}$, a contradiction.
[…]
(a) The case $|X\cap Y|=1$.
$x_{0}^{\prime}$$x_{0}$$x_{1}$$_{0}yy$$y_{4}$$a_{1}$$a_{4}$$a_{2}$$a_{3}$
(b) The case $|X\cap Y|=0$.
Figure 4: The case $|X\cap C_{0}|=2$ in the proof of Lemma 3.2. Here a dashed line indicates that the two vertices may or may not be adjacent.
If $|X\cap C_{0}|=1$, consider a vertex set $Y\subseteq C_{0}\cup C_{1}\cup C_{4}$ that induces a $K_{3}$. If $|Y\cap C_{0}|=2$, then by the arguments in the preceding paragraphs, we deduce that $C_{0}\cup C_{2}\cup C_{3}$ is $K_{3}$-free, a contradiction. Hence, $|Y\cap C_{0}|=1$. Let $X=\{x_{0},x_{1},x_{2}\}$ and $Y=\{y_{0},y_{1},y_{4}\}$, where $x_{0},y_{0}\in C_{0}$, $x_{1},y_{1}\in C_{1}$, $x_{2}\in C_{2}$, and $y_{4}\in C_{4}$. Since $k=3$ and $G$ is diamond-free, we have $|X\cap Y|\leq 1$.
For the case $|X\cap Y|=1$, either $x_{0}=y_{0}$ or $x_{1}=y_{1}$. If $x_{0}=y_{0}$ (see Figure 5(a)), then $\{x_{1},x_{2}\}$ is anticomplete to $\{y_{1},y_{4}\}$ since $H$ is $K_{4}$-free and diamond-free. Thus, $\{x_{1},x_{2},y_{1},y_{4},a_{4},a_{3}\}$ induces a $P_{2}\cup P_{4}$, a contradiction. If $x_{1}=y_{1}$ (see Figure 5(b)), then $\{x_{0},x_{2}\}$ is anticomplete to $\{y_{0},y_{4}\}$ since $H$ is $K_{4}$-free and diamond-free. Thus, $\{x_{0},x_{2},y_{0},y_{4},a_{4},a_{3}\}$ induces a $P_{2}\cup P_{4}$, a contradiction. For the case $|X\cap Y|=0$ (see Figure 5(c)), since $k=3$ and $G$ is diamond-free, $\{x_{0},x_{1}\},Y]$ is a matching in $G$. If $\{x_{0},x_{1}\}$ is anticomplete to $Y$, then $\{x_{0},x_{1},y_{0},y_{4},a_{4},a_{3}\}$ induces a $P_{2}\cup P_{4}$, a contradiction. Hence, $\{x_{0},x_{1}\},Y]$ is a nonempty matching in $G$. By Lemma 1.4, $G[\{x_{0},x_{1}\}\cup Y]$ contains a $P_{4}$. The vertices of this $P_{4}$, together with $\{a_{2},a_{3}\}$, induce a $P_{2}\cup P_{4}$, a contradiction.
$x_{2}$$x_{0}=y_{0}$$y_{4}$$x_{1}$$y_{1}$$a_{1}$$a_{2}$$a_{3}$$a_{4}$
(a) The case $x_{0}=y_{0}$.
$x_{2}$$x_{1}=y_{1}$$y_{0}$$x_{0}$$y_{4}$$a_{1}$$a_{2}$$a_{3}$$a_{4}$
(b) The case $x_{1}=y_{1}$.
$x_{2}$$x_{0}$$y_{0}$$y_{4}$$x_{1}$$y_{1}$$a_{1}$$a_{2}$$a_{3}$$a_{4}$
(c) The case $|X\cap Y|=0$.
Figure 5: The case $|X\cap C_{0}|=1$ in the proof of Lemma 3.2.
Thus, at least one of $G[C_{0}\cup C_{1}\cup C_{2}]$, $G[C_{0}\cup C_{3}\cup C_{4}]$, and $G[C_{0}\cup C_{1}\cup C_{4}]$ is $K_{3}$-free, and hence $\chi(G)=4$. ∎
Proof of Lemma 3.3.
We first show that if there exist three sets among $C_{1},\dots,C_{5}$ such that the subgraph induced by their union is $K_{3}$-free, then $\chi(G)=5$. Without loss of generality, suppose that $G[C_{1}\cup C_{2}\cup C_{3}]$ is $K_{3}$-free. Then $\omega(C_{1}\cup C_{2}\cup C_{3})\leq 2$. By Lemma 2.1, both $C_{1}\cup C_{2}\cup C_{3}$ and $C_{0}\cup C_{4}\cup C_{5}$ induce perfect graphs. Thus, we may colour $a_{i}$ with $i$ for $i\in[5]$, colour $C_{1}\cup C_{2}\cup C_{3}$ with colours from $\{4,5\}$, and colour $C_{0}\cup C_{4}\cup C_{5}$ with colours from $\{1,2,3\}$.
Hence, we may assume that for any three sets among $C_{1},\dots,C_{5}$, the subgraph induced by their union contains a $K_{3}$. Since $k=3$, we may regard $B$ as such a $K_{3}$. Then, by ($\ast$2), each such $K_{3}$ has its three vertices belonging to three different sets among $C_{1},\dots,C_{5}$. Let $X=\{x_{1},x_{2},x_{3}\}\subseteq C_{1}\cup C_{2}\cup C_{3}$ and $Y=\{y_{1},y_{2},y_{3}\}\subseteq C_{1}\cup C_{4}\cup C_{5}$ be two vertex sets such that both $X$ and $Y$ induce a $K_{3}$, where $x_{i}\in C_{i}$ for $i\in\{1,2,3\}$ and $y_{j}\in C_{j}$ for $j\in\{1,4,5\}$.
If $x_{1}=y_{1}$ (see Figure 6(a)), since $k\leq 3$ and $G$ is diamond-free, $\{x_{2},x_{3}\}$ is anticomplete to $\{y_{4},y_{5}\}$. Hence, $\{x_{2},x_{3},y_{4},y_{5},a_{5},a_{1}\}$ induces a $P_{2}\cup P_{4}$, a contradiction.
If $x_{1}\neq y_{1}$ (see Figure 6(b)), since $k\leq 3$ and $G$ is diamond-free, $\{x_{1},x_{2}\},Y]$ is a matching in $G$. If $\{x_{1},x_{2}\}$ is anticomplete to $Y$, then $\{x_{1},x_{2},y_{1},y_{5},a_{5},a_{4}\}$ induces a $P_{2}\cup P_{4}$, a contradiction. Hence, the matching $\{x_{1},x_{2}\},Y]$ is nonempty. By Lemma 1.4, $G[\{x_{1},x_{2}\}\cup Y]$ contains a $P_{4}$ whose vertex set contains $y_{1}$. Thus, $\{x_{1},x_{2},y_{1},y_{4}\}$ or $\{x_{1},x_{2},y_{1},y_{5}\}$ induces a $P_{4}$; hence, $\{a_{3},a_{5},x_{1},x_{2},y_{1},y_{4}\}$ or $\{a_{3},a_{4},x_{1},x_{2},y_{1},y_{5}\}$ induces a $P_{2}\cup P_{4}$. In either case, we obtain a contradiction.
$x_{3}$$x_{1}=y_{1}$$y_{4}$$y_{5}$$x_{2}$$a_{1}$$a_{2}$$a_{3}$$a_{4}$$a_{5}$
(a) The case $x_{1}=y_{1}$.
$x_{3}$$x_{1}$$y_{1}$$y_{4}$$y_{5}$$x_{2}$$a_{1}$$a_{2}$$a_{3}$$a_{4}$$a_{5}$
(b) The case $x_{1}\neq y_{1}$.
Figure 6: Two cases in the proof of Lemma 3.3.
Thus, at least one of $G[C_{1}\cup C_{2}\cup C_{3}]$ and $G[C_{1}\cup C_{4}\cup C_{5}]$ is $K_{3}$-free, and hence $\chi(G)=5$. ∎
4 The Case $k\geq 4$
In this case, $\omega\geq k\geq 4$. This section aims to prove the following lemma.
Lemma 4.1.
If $G$ is a ($P_{2}\cup P_{4}$, diamond)-free graph with $\omega\geq 4$ and $k\geq 4$, then $\chi(G)=\omega$.
To prove Lemma 4.1, we first present the following claim.
Claim 4.1.
Let $x,y$ be two vertices in $R\cup S\cup T\cup Z$. If $N_{A\cup B}(x)\neq N_{A\cup B}(y)$, and $|N_{A\cup B}(x)\cup N_{A\cup B}(y)|\leq 3$, then $x\nsim y$.
Proof.
Suppose, to the contrary, that there exist vertices $x,y\in R\cup S\cup T\cup Z$ satisfying the above conditions such that $x\sim y$. Since each of $|N_{A}(x)|$, $|N_{A}(y)|$, $|N_{B}(x)|$, and $|N_{B}(y)|$ is at most $1$, we may discuss the following two cases separately.
Case 1. $|N_{A}(x)\cup N_{A}(y)|=2$ or $|N_{B}(x)\cup N_{B}(y)|=2$.
Suppose $|N_{A}(x)\cup N_{A}(y)|=2$ (the case where $|N_{B}(x)\cup N_{B}(y)|=2$ is analogous). Since $|N_{A\cup B}(x)\cup N_{A\cup B}(y)|\leq 3$, we have $|N_{B}(x)\cup N_{B}(y)|\leq 1$. By symmetry, we may assume that $x\in C_{1}\cap(C_{0}^{\prime}\cup C_{1}^{\prime})$ and $y\in C_{2}\cap(C_{0}^{\prime}\cup C_{1}^{\prime})$. Hence, $\{x,y\}$ is anticomplete to $\{a_{3},\ldots,a_{\omega},b_{2},\ldots,b_{k}\}$. By ($\ast$2), $\{a_{3},\ldots,a_{\omega}\},\{b_{2},\ldots,b_{k}\}]$ is a matching. Since $\omega,k\geq 4$ and $G$ is $(P_{2}\cup P_{4})$-free, Lemma 1.4 implies that $\{a_{3},\ldots,a_{\omega}\}$ is anticomplete to $\{b_{2},\ldots,b_{k}\}$. Since $k\geq 4$, there exist vertices $b_{i},b_{j}\in\{b_{2},\ldots,b_{k}\}$ that are anticomplete to $a_{1}$. Consequently, $\{b_{i},b_{j},a_{3},a_{1},x,y\}$ induces a $P_{2}\cup P_{4}$, a contradiction.
Case 2. $|N_{A}(x)\cup N_{A}(y)|\leq 1$ and $|N_{B}(x)\cup N_{B}(y)|\leq 1$.
By symmetry, we may assume that $x,y\in(C_{0}\cup C_{1})\cap(C_{0}^{\prime}\cup C_{1}^{\prime})$. Thus, $\{x,y\}$ is anticomplete to $\{a_{2},\ldots,a_{\omega},b_{2},\ldots,b_{k}\}$. By ($\ast$2), $\{a_{2},\ldots,a_{\omega}\},\{b_{2},\ldots,b_{k}\}]$ is a matching. Since $\omega,k\geq 4$ and $G$ is $(P_{2}\cup P_{4})$-free, Lemma 1.4 implies that $\{a_{2},\ldots,a_{\omega}\}$ is anticomplete to $\{b_{2},\ldots,b_{k}\}$. Since $N_{A\cup B}(x)\neq N_{A\cup B}(y)$, $N_{A}(x)\neq N_{A}(y)$ or $N_{B}(x)\neq N_{B}(y)$. Suppose $N_{A}(x)\neq N_{A}(y)$ (the case where $N_{B}(x)\neq N_{B}(y)$ is analogous). Then either $a_{1}\sim x$ and $a_{1}\nsim y$, or $a_{1}\nsim x$ and $a_{1}\sim y$. By symmetry, assume that $a_{1}\sim x$ and $a_{1}\nsim y$. Since $k\geq 4$, there exist vertices $b_{i},b_{j}\in\{b_{2},\ldots,b_{k}\}$ that are anticomplete to $a_{1}$. Consequently, $\{b_{i},b_{j},a_{2},a_{1},x,y\}$ induces a $P_{2}\cup P_{4}$, a contradiction.
This completes the proof of Claim 4.1
Then we present several useful properties of the subsets in the partition defined in Section 2. In particular, Properties (P1)–(P4) follow directly from Claim 4.1.
(P1) $Z$ is anticomplete to $R\cup S\cup T$.
(P2) $R$ is anticomplete to $S\cup T\cup Z$, and $R_{i}$ is anticomplete to $R_{j}$ for distinct $i,j\in[k]$.
(P3) $T$ is anticomplete to $R\cup S\cup Z$, and $T_{i}$ is anticomplete to $T_{j}$ for distinct $i,j\in[\omega]$.
(P4) $S$ is anticomplete to $R\cup T\cup Z$. Moreover, $S_{i}^{p}$ is anticomplete to $S_{i}^{q}$ for every $i\in[\omega]$ and distinct $p,q\in[k]$; and $S_{i}^{p}$ is anticomplete to $S_{j}^{p}$ for every $p\in[k]$ and distinct $i,j\in[\omega]$.
(P5) Suppose $\omega=4$. If $S_{i}^{p}$ is not anticomplete to $S_{j}^{q}$, then $S_{i}^{p}$ is anticomplete to $S_{j}^{q^{\prime}}$ and $S_{j^{\prime}}^{q}$ for pairwise distinct $p,q,q^{\prime}\in[4]$ and pairwise distinct $i,j,j^{\prime}\in[4]$.
Proof.
By symmetry, it suffices to show that if $S_{1}^{1}$ is not anticomplete to $S_{2}^{2}$, then $S_{1}^{1}$ is anticomplete to $S_{2}^{3}$ and $S_{3}^{2}$. Let $x\in S_{1}^{1}$ and $y\in S_{2}^{2}$ with $x\sim y$.
If $b_{3}$ is anticomplete to $\{a_{3},a_{4}\}$, then $|[\{b_{3},b_{4}\},\{a_{1},a_{2}\}]|=2$ and $|[\{a_{3},a_{4}\},\{b_{1},b_{2}\}]|=2$. Otherwise, if $|[\{b_{3},b_{4}\},\{a_{1},a_{2}\}]|\leq 1$, then there exist $a_{i}\in\{a_{1},a_{2}\}$ and $a_{j}\in\{a_{3},a_{4}\}$ such that both $a_{i}$ and $a_{j}$ are anticomplete to $\{b_{3},b_{4}\}$, and hence $\{b_{3},b_{4},a_{j},a_{i},x,y\}$ induces a $P_{2}\cup P_{4}$, a contradiction. Similarly, if $|[\{a_{3},a_{4}\},\{b_{1},b_{2}\}]|\leq 1$, then there exists $b_{p}\in\{b_{1},b_{2}\}$ anticomplete to $\{a_{3},a_{4}\}$, and hence $\{a_{3},a_{4},b_{3},b_{p},x,y\}$ induces a $P_{2}\cup P_{4}$, a contradiction.
If $b_{3}$ is not anticomplete to $\{a_{3},a_{4}\}$, then since $\{x,y,b_{3},b_{4},a_{3},a_{4}\}$ cannot induce a $P_{2}\cup P_{4}$, it follows that $|[\{b_{3},b_{4}\},\{a_{3},a_{4}\}]|=2$.
We now show that $S_{1}^{1}$ is anticomplete to $S_{2}^{3}$; the argument for $S_{1}^{1}$ being anticomplete to $S_{3}^{2}$ is analogous. Suppose, to the contrary, that there exists a vertex $z\in S_{2}^{3}$ such that $x\sim z$. If $|[\{b_{3},b_{4}\},\{a_{1},a_{2}\}]|=2$ and $|[\{a_{3},a_{4}\},\{b_{1},b_{2}\}]|=2$, then $\{a_{3},a_{4}\}$ is anticomplete to $b_{4}$ and $\{a_{1},a_{2}\}$ is anticomplete to $b_{2}$. Moreover, there exist vertices $u\in\{a_{3},a_{4}\}$ nonadjacent to $b_{2}$ and $v\in\{a_{1},a_{2}\}$ nonadjacent to $b_{4}$. Thus, $\{b_{2},b_{4},u,v,x,z\}$ induces a $P_{2}\cup P_{4}$, a contradiction. If $|[\{b_{3},b_{4}\},\{a_{3},a_{4}\}]|=2$, then $\{a_{1},a_{2}\}$ is anticomplete to $b_{4}$ and $\{a_{3},a_{4}\}$ is anticomplete to $b_{2}$. Moreover, there exist vertices $u\in\{a_{3},a_{4}\}$ nonadjacent to $b_{4}$ and $v\in\{a_{1},a_{2}\}$ nonadjacent to $b_{2}$. Therefore, $\{b_{2},b_{4},u,v,x,z\}$ induces a $P_{2}\cup P_{4}$, a contradiction. ∎
(P6) If $\omega=4$, then the vertex set of each component in $G[S]$ can be partitioned into at most four stable sets, each contained in a distinct $S_{i}^{j}$, with all $i$’s and all $j$’s being distinct within that component.
Proof.
This follows directly from (P4) and (P5). ∎
(P7) If $\omega\geq 5$, then $S_{i}^{p}$ is anticomplete to $S_{j}^{q}$ for distinct $p,q\in[k]$ and distinct $i,j\in[\omega]$.
Proof.
By symmetry, it suffices to show that $S_{1}^{1}$ is anticomplete to $S_{2}^{2}$. Suppose to the contrary that there exist vertices $x\in S_{1}^{1}$ and $y\in S_{2}^{2}$ such that $x\sim y$. By ($\ast$2), $\{a_{3},\ldots,a_{\omega}\},\{b_{3},\ldots,b_{k}\}]$ is a matching. Since $\omega\geq 5$, $k\geq 4$ and $G$ is $(P_{2}\cup P_{4})$-free, Lemma 1.4 implies that $\{a_{3},\ldots,a_{\omega}\}$ is anticomplete to $\{b_{3},\ldots,b_{k}\}$. Since $\omega\geq 5$, there exist vertices $a_{i},a_{j}\in\{a_{3},\ldots,a_{\omega}\}$ that are anticomplete to $b_{1}$. Consequently, $\{a_{i},a_{j},b_{3},b_{1},x,y\}$ induces a $P_{2}\cup P_{4}$, a contradiction. ∎
With these properties in hand, we now prove Lemma 4.1.
Proof of Lemma 4.1.
Let $G$ be a $(P_{2}\cup P_{4},\text{ diamond})$-free graph with $\omega\geq 4$ and $k\geq 4$. By ($\ast\ast$2), $A$ is not anticomplete to $B$. Then, by ($\ast$2), we may relabel $B$ so that, for some integer $m\in[k]$, we have $a_{i}\sim b_{i}$ for all $i\in[m]$, and $a_{p}\nsim b_{q}$ whenever $p\in[\omega]\setminus[m]$ and $q\in[k]\setminus[m]$. By Lemma 2.1, each of the sets $R_{j}$, $T_{i}$, and $Z$ induces a perfect graph. Since $\omega(R_{j}),\omega(T_{i})\leq\omega-1$ and $\omega(Z)\leq k\leq\omega$, we may colour them with at most $\omega-1$, $\omega-1$, and $\omega$ colours, respectively. Now we construct an $\omega$-colouring of $G$.
We first colour $A$ and $B$.
  • Colour $a_{i}$ with colour $i$ for $i\in[\omega]$.
  • if $m=1$, colour $b_{i}$ with colour $i+1$ for $i\in[k-1]$, and colour $b_{k}$ with colour $1$;
  • if $m\geq 2$, colour $b_{i}$ with colour $i+1$ for $i\in[m-1]$, colour $b_{m}$ with colour $1$, and colour $b_{j}$ with colour $j$ for $j\in[k]\setminus[m]$.
Next, we colour $S$, considering separately the cases $\omega=4$ and $\omega\geq 5$.
  • If $\omega=4$, by (P6), we may colour each component of $G[S]$ with at most four suitable colours from $[4]$. To justify this, we construct a bipartite graph demonstrating the existence of such a colouring. Let $X$ denote the set of stable sets in a component of $G[S]$ as described in (P6). For each $X_{t}\in X$, let $Y_{t}\subseteq[4]$ be the set of colours that can be assigned to $X_{t}$; for example, if $X_{t}\subseteq S_{i}^{j}$, then $Y_{t}=[4]\setminus(c(a_{i})\cup c(b_{j}))$. Let
    $Y=\bigcup_{X_{t}\in X}Y_{t}.$
    We now regard $X$ and $Y$ as the two vertex sets of a bipartite graph $Q$ with bipartition $(X,Y)$. Each vertex $X_{t}\in X$ represents a stable set in the component, and each vertex $y\in Y$ represents a colour in $[4]$. Vertices $X_{t}$ and $y$ are adjacent in $Q$ if and only if $y\in Y_{t}$; equivalently, $N_{Q}(X_{t})=Y_{t}$ for all $X_{t}\in X$. By Hall’s Theorem [9] applied to $Q$, there exists a matching saturating $X$.
  • If $\omega\geq 5$, by (P4) and (P7), we may colour each $S_{i}^{j}$ with one colour from $\omega]\setminus(c(A_{i})\cup c(B_{j}))$ for $i\in[\omega]$ and $j\in[k]$.
Finally, we colour $Z$, $R$, and $T$ as follows:
  • By (P1), colour $Z$ with at most $\omega$ colours from $\omega]$.
  • By (P2), colour $R_{i}$ with at most $\omega-1$ colours from $\omega]\setminus c(B_{i})$ for $i\in[k]$.
  • By (P3), colour $T_{i}$ with at most $\omega-1$ colours from $\omega]\setminus c(A_{i})$ for $i\in[\omega]$.
This completes the proof of Lemma 4.1. ∎
5 Proof of Theorem 1.3
(a) The Mycielski–Grötzsch graph.
(b) The complement of the Schläfli graph.
Figure 7: The graphs $H_{1}$ and $H_{2}$ used to show the optimality of the bounds in Theorem 1.3.
Proof of Theorem 1.3.
Let $G$ be a ($P_{2}\cup P_{4}$, diamond)-free graph. If $\omega\leq 2$, then $\chi(G)\leq 4$ by Lemma 2.2; thus we may assume $\omega\geq 3$. If $k\leq 2$, then the theorem holds by Lemmas 2.3 and 2.4; hence we may further assume $k\geq 3$.
If $\omega=3$ and $k=3$, then Lemma 3.1 implies $\chi(G)\leq 6$.
If $\omega\geq 4$ and $k=3$, then Lemmas 2.3, 3.2, and 3.3 imply $\chi(G)=\omega$.
If $\omega\geq 4$ and $k\geq 4$, then Lemma 4.1 implies $\chi(G)=\omega$.
Next, we show that the bounds are optimal for all $\omega\geq 2$ by presenting two examples for $\omega=2$ and $\omega=3$ that attain the corresponding upper bounds.
Let $H_{1}$ be the Mycielski–Grötzsch graph (see Figure 7(a)) and let $H_{2}$ be the complement of the Schläfli graph (see Figure 7(b); also available at https://houseofgraphs.org/graphs/19273). Both $H_{1}$ and $H_{2}$ are $(P_{2}\cup P_{4},\text{ diamond})$-free. Moreover, $\omega(H_{1})=2$ and $\chi(H_{1})=4$, while $\omega(H_{2})=3$ and $\chi(H_{2})=6$.
This completes the proof of Theorem 1.3. ∎
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
No data was used for the research described in the article.